OpenStudy (anonymous):

If a,b are the roots of equation x^2+3x+4=0 then a+b = ?

OpenStudy (anonymous):

That equation has no real roots. But actually has complex ones and they can be added Symply by using the following formula: If x1 and x2 are the roots of the ecuation ax^2+bx+c =0 Then x1+x2 = -b/a In this case that happens to be -3.